# Speedy-Splat: Fast 3D Gaussian Splatting with Sparse Pixels and Sparse Primitives

### v1

@@ Abstract
3D Gaussian Splatting (3D-GS) is a recent 3D scene reconstruction technique that enables real-time rendering of novel views by modeling scenes as parametric point clouds of differentiable 3D Gaussians. However, its rendering speed and model size still present bottlenecks, especially in resource-constrained settings. In this paper, we identify and address two key inefficiencies in 3D-GS, achieving substantial improvements in rendering speed, model size, and training time. First, we optimize the rendering pipeline to precisely localize Gaussians in the scene, boosting rendering speed without altering visual fidelity. Second, we introduce a novel pruning technique and integrate it into the training pipeline, significantly reducing model size and training time while further raising rendering speed. Our Speedy-Splat approach combines these techniques to accelerate average rendering speed by a drastic $\mathit{6.71\times}$ across scenes from the Mip-NeRF 360, Tanks & Temples, and Deep Blending datasets with $\mathit{10.6\times}$ fewer primitives than 3D-GS.
@@ 1 Introduction
Fast rendering of photorealistic novel views has been a long-standing goal in computer vision and graphics. Neural Radiance Fields (NeRF) [20] and its variants have made significant strides in photorealistic 3D scene reconstruction by representing scenes as continuous neural volumetric models that encode scene density and color at spatial coordinates. However, despite recent efforts [21, 4, 24], fast rendering in NeRF remains a challenge, as the the volumetric sampling used in ray-marching is computationally expensive. Recently, 3D Gaussian Splatting (3D-GS) [14] has emerged as a promising alternative that enables real-time rendering by modeling scenes as parametric point clouds of differentiable 3D Gaussians. Nevertheless, its rendering speed is still limited in practice by high parameter counts and certain algorithmic inefficiencies. Efficient rendering is essential in settings like virtual reality, networked applications, and multi-view streaming, and real-time rendering on resource-constrained edge devices like mobile phones is yet to be achieved [18].
@@
We begin by observing that the cost of 3D-GS rendering is proportional to both the number of Gaussians in the scene and the number of pixels processed per Gaussian. Our approach optimizes both factors. First, we reduce the number of pixels that are processed for each Gaussian by accurately localizing it in the rendered image. Second, we reduce the total number of Gaussians in the model through a novel approach that maintains rendering quality.
3D-GS implements tile rendering to localize Gaussians in the image plane, assigning them to only the image tiles that they intersect. We find the existing algorithm to be overly conservative; in Section 4.1.1, we address this by introducing our SnugBox algorithm to precisely localize Gaussians by computing a tight bounding box around their extent. Then, in Section  4.1.2, we extend SnugBox with our AccuTile algorithm to identify exact tile intersections. Both approaches are plug-and-play, lead to respective inference speed-ups of 1.82×1.82$\times$1.82 × and 1.99×1.$9$ on average, and do not change the 3D-GS renderings.
To reduce model size while preserving visual fidelity, we reduce the memory requirement of an an existing pruning method [11] by 36×36$\times$36 × and simplify its integration into the 3D-GS training pipeline. Gaussians are replicated and pruned at regular intervals during the densification stage of 3D-GS training. In Section 4.2.2, we augment the densification stage with our Soft Pruning method to prune $80\%$ of Gaussians at regular intervals. After the densification stage is completed, we use our Hard Pruning method, described in Section 4.2.3, to prune an additional $30\%$ of Gaussians at set intervals.
@@ 2 Related work
The real-time rendering speed of 3D-GS [14] on desktop GPUs has sparked research interest in further accelerating its training and inference in resource-constrained environments. In this section, we discuss several related works that have emerged in the literature.
@@ 2.2 Other Methods
In addition to pruning, several other strategies have been explored to enhance the rendering and training speed of 3D-GS. Mini-Splatting [7] modifies its densification strategies and adds a simplification stage to constrain the number of Gaussians. 3DGS-LM [13] replaces Adam [15] with a tailored Levenberg-Marquardt optimizer [9]. Taming 3DGS [25] proposes a constructive optimization process that limits the number of Gaussians to a pre-defined threshold set by the user. DISTWAR [5] dives into the low-level implementation of GPU thread scheduling and optimizes atomic processing with a novel primitive. FlashGS [8] is a recent concurrent work that also uses efficient tile intersection to improve rendering speed; however, their overall speedup is lower than ours because they do not explore pruning. Most of these approaches are orthogonal to ours and can be applied alongside it.
@@ 3 Background
Table 1: Average execution time (milliseconds) of each function across all scenes. The operation in each row is applied cumulatively to all of the following rows. For each model, accurate measurements are collected by averaging execution times across three runs that each render the test set 20 times to reduce variance. The fastest and second fastest times are color coded.
@@ 3.2.1 Preprocessing
where $\boldsymbol{J}$ is the Jacobian of the first order approximation of the perspective projection. Dropping the last row and column of $\boldsymbol{\hat{\Sigma}_{i_{3D}}}$ gives $\boldsymbol{\Sigma_{i_{2D}}}$ [14, 26]. The largest eigenvalue of $\boldsymbol{\Sigma_{i_{2D}}}$ is used to compute the count of tiles intersected by this Gaussian $\mathcal{G}_{i}$ as shown in Figure 2(a). $\boldsymbol{\Sigma_{i_{2D}}}^{-1}$ is computed, and a view dependent color $c_{i}$ is derived from $W$ and $h_{i}$. All three values are stored for later processing.
@@ 4.1 Precise Tile Intersect
$r=\lceil 3\sqrt{\lambda_{\max}}\rceil.$ (8)
@@ 4.1.2 AccuTile
Algorithm 1 The AccuTile Algorithm. For simplicity, the algorithm outlined here is applied to the rows of the SnugBox tile extent bounding box, matching the example in Figure 2(c). In practice, it is applied along the smaller side of the tile extent. The subscripts $t$, $b$, $l$, and $r$ represent the top, bottom, left, and right sides, respectively.
@@ 4.2 Efficient Pruning
Although PUP 3D-GS touts high compression ratios and rendering speeds, we identify two key drawbacks in its formulation. First, computing the Hessian requires storage proportional to $N\times 36$, where $N$ is the number of Gaussians, when the memory footprint of the 3D-GS model, which has $59$ parameters per Gaussian, is proportional to $N\times 59$. While their score works well for post-hoc pruning, it incurs a resource constraint for use during training.
@@
Our approach builds on PUP 3D-GS by introducing an efficient pruning score that we integrate into the 3D-GS training pipeline. Then, we define two distinct modalities for pruning the model during training: Soft Pruning occurs during densification in the first $15000$ iterations, and Hard Pruning is performed after densification is completed after iteration $15000$.
@@ 4.2.3 Hard Pruning
We also observe that the model performance after densification closely matches that of the fully-trained model. The iterations after densification is completed essentially fine-tune the model – we can use them to ”refine” the models after pruning like PUP 3D-GS [11] and LightGaussian [6]. In practice, our Hard Pruning method prunes the model by a constant ratio every $3000$ iterations starting at iteration $15000$. When paired with soft pruning, we Hard Prune $30\%$ of Gaussians in each interval which empirically reduces the total number of Gaussians across scenes by 10.6×10.6$\times$10.6 ×.
@@ 5.2 Implementation Details
Table 2: Average Gaussian count, FPS, and training time across all scenes. Ratios for model size compression, rendering speed-up, and training speed-up are reported in (parentheses). The operation in each row is applied cumulatively to all of the following rows. The best and second best value for each metric are color coded.
@@
Table 3: Average reported metrics for each pruning method across all scenes in the Mip-NeRF 360 dataset. The Comp column reports model size compression in terms of Gaussian count, FPS reports rendering speed-up, and Train reports training time speed-up, all with respect to the baseline 3D-GS model. PSNR, SSIM, and LPIPS are also recorded. The best and second best value for each metric are color coded; lossless methods are underlined. Results for the Tanks & Temples and Deep Blending datasets are reported in Appendix A.1.
@@ 5.3.2 Overall Performance
Table 3 compares our methods against other methods that reduce the number of Gaussians and increase inference speed using the mean of each metric across all scenes in the Mip-NeRF 360 dataset. The underlined methods are ”lossless”, meaning that they avoid degrading visual fidelity at all. SnugBox and AccuTile, our lossless methods, improve rendering and training speed while leaving image quality metrics essentially unchanged or slightly better. Our full pipeline, labeled as “+Hard Pruning” boasts the highest compression ratios, rendering speeds, and training speed-ups across all datasets. Furthermore, its image quality metrics are competitive with the other methods.
@@ 5.4 Pruning Score Comparison
Table 4: Average metrics across all scenes when using the Speedy-Splat and PUP 3D-GS pruning scores to prune $88.44\%$ of Gaussians using the PUP 3D-GS pipeline. Two rounds of prune-refine are performed on each baseline 3D-GS model, pruning $66\%$ of Gaussians and then fine-tuning for $5,000$ iterations in each one. The best value for each metric is color coded.
@@ Appendix A Appendix
A.1 Additional Datasets Evaluation
@@
A.2 Per-Scene Metrics
PSNR, SSIM, LPIPS, FPS, and training times for each scene from the Mip-NeRF 360, Tanks&Temples, and Deep Blending datasets that was used in 3D-GS [14] are recorded in Tables 7, 8, 9,  10, and Table 11, respectively. The operation in each row is applied cumulatively to all of the following rows.

### v2

@@ Abstract
3D Gaussian Splatting (3D-GS) is a recent 3D scene reconstruction technique that enables real-time rendering of novel views by modeling scenes as parametric point clouds of differentiable 3D Gaussians. However, its rendering speed and model size still present bottlenecks, especially in resource-constrained settings. In this paper, we identify and address two key inefficiencies in 3D-GS to substantially improve rendering speed. These improvements also yield the ancillary benefits of reduced model size and training time. First, we optimize the rendering pipeline to precisely localize Gaussians in the scene, boosting rendering speed without altering visual fidelity. Second, we introduce a novel pruning technique and integrate it into the training pipeline, significantly reducing model size and training time while further raising rendering speed. Our Speedy-Splat approach combines these techniques to accelerate average rendering speed by a drastic $\mathit{6.71\times}$ across scenes from the Mip-NeRF 360, Tanks & Temples, and Deep Blending datasets.
@@ 1 Introduction
Fast rendering of photorealistic novel views has been a long-standing goal in computer vision and graphics. Neural Radiance Fields (NeRF) [23] and its variants have made significant strides in photorealistic 3D scene reconstruction by representing scenes as continuous neural volumetric models that encode scene density and color at spatial coordinates. However, despite recent efforts [24, 4, 29], fast rendering in NeRF remains challenging because the volumetric sampling used in ray-marching is computationally expensive. Recently, 3D Gaussian Splatting (3D-GS) [14] has emerged as a promising alternative that enables real-time rendering by modeling scenes as parametric point clouds of differentiable 3D Gaussians. Nevertheless, its rendering speed is still limited by high parameter counts and certain algorithmic inefficiencies. Efficient rendering is essential in applications such as virtual reality, networked systems, and multi-view streaming. Real-time rendering on resource-constrained edge devices, such as mobile phones, has yet to be achieved [19].
@@
We begin by observing that the cost of 3D-GS rendering is proportional to both the number of Gaussians in the scene and the number of pixels processed per Gaussian. Our approach optimizes both factors. First, we reduce the number of pixels that are processed for each Gaussian by efficiently and accurately localizing it in the rendered image. Second, we reduce the total number of Gaussians in the model through a novel approach that maintains rendering quality.
3D-GS implements tile rendering to localize Gaussians in the image plane by assigning each Gaussian to the tiles that it intersects. We find the existing algorithm to be overly conservative; in Section 4.1.1, we address this by introducing our SnugBox algorithm to precisely localize Gaussians by computing a tight bounding box around their extent. Then, in Section  4.1.2, we extend SnugBox with our AccuTile algorithm to identify exact tile intersections. Both approaches are plug-and-play, lead to respective inference speed-ups of 1.82×1.82$\times$1.82 × and 1.99×1.$9$ on average, and do not change the 3D-GS renderings.
To reduce the total number of Gaussians while preserving visual fidelity, we extend an existing pruning method [11] by reducing its memory requirement by 36×36$\times$36 × and incorporating it into the 3D-GS training pipeline. During the densification stage of 3D-GS training, Gaussians are regularly replicated and pruned. In Section 4.2.2, we augment the densification stage with our Soft Pruning method to prune $80\%$ of Gaussians at regular intervals. After the densification stage, our Hard Pruning method, described in Section 4.2.3, prunes an additional $30\%$ of Gaussians at set intervals.
@@ 2 Related work
The real-time rendering speed of 3D-GS [14] on desktop GPUs has inspired research focused on further accelerating both its training and inference in resource-constrained environments. In this section, we review related works that specifically target these performance improvements.
@@ 2.2 Other Methods
In addition to pruning, several other strategies have been explored to enhance the rendering and training speed of 3D-GS. Mini-Splatting [7] modifies its densification strategies and adds a simplification stage to constrain the number of Gaussians. 3DGS-MCMC [15] models training dynamics as an MCMC process. Revisiting Densification [30] introduces an error-based densification strategy. 3DGS-LM [13] replaces Adam [16] with a tailored Levenberg-Marquardt optimizer [9]. Taming 3DGS [22] proposes a constructive optimization process that limits the number of Gaussians to a pre-defined threshold set by the user. DISTWAR [5] dives into the low-level implementation of GPU thread scheduling and optimizes atomic processing with a novel primitive. StopThePop [27] introduces a precise tile intersection method, but we find that our approach is notably faster in Appendix A.4. FlashGS [8] implements a tile intersection method similar to StopThePop. Several works reduce training time and memory requirements by enforcing geometric constraints [21, 28, 32, 31]. Most of these approaches are orthogonal to ours and can be applied alongside it.
@@ 3 Background
Table 1: Average execution time (milliseconds) of each function across all scenes in Section 5.1. The operation in each row is applied cumulatively to all of the following rows. For each model, accurate measurements are collected by averaging execution times across three runs that each render the test set 20 times to reduce variance. The fastest and second fastest times are color coded.
@@ 3.2.1 Preprocessing
where $\boldsymbol{J}$ is the Jacobian of the first order approximation of the perspective projection. Dropping the last row and column of $\boldsymbol{\hat{\Sigma}_{i_{3D}}}$ gives $\boldsymbol{\Sigma_{i_{2D}}}$ [14, 33]. The largest eigenvalue of $\boldsymbol{\Sigma_{i_{2D}}}$ is used to compute the count of tiles intersected by this Gaussian $\mathcal{G}_{i}$ as shown in Figure 2(a). $\boldsymbol{\Sigma_{i_{2D}}}^{-1}$ is computed, along with a view-dependent color $c_{i}$, derived from $W$ and $h_{i}$. All three values are stored for later processing.
@@ 4.1 Precise Tile Intersect
$r=\left\lceil 3\sqrt{\lambda_{\max}}\right\rceil.$ (8)
@@ 4.1.2 AccuTile
Algorithm 1 The AccuTile Algorithm. For simplicity, the algorithm outlined here is applied to the rows of the SnugBox tile extent bounding box, matching the example in Figure 2(c). In practice, it is applied along the smaller side of the tile extent. The subscripts $t$, $b$, $l$, and $r$ represent the top, bottom, left, and right sides, respectively. A proof of correctness sketch is presented in Appendix A.1.
@@ 4.2 Efficient Pruning
Although PUP 3D-GS touts high compression ratios and rendering speeds, we identify two key drawbacks in its formulation. First, computing the Hessian requires storage proportional to $N\times 36$, where $N$ is the number of Gaussians. In comparison, the 3D-GS model has a memory footprint proportional to $N\times 59$ because it stores $59$ parameters per Gaussian. While this score is effective for post-hoc pruning, using it during training is impractical.
@@
Our approach builds on PUP 3D-GS by introducing an efficient pruning score that we incorporate into the 3D-GS training pipeline. We also define two distinct pruning modalities: Soft Pruning, which takes place during densification in the first $15000$ iterations, and Hard Pruning, which is applied after densification is completed after iteration $15000$.
@@ 4.2.3 Hard Pruning
We also observe that the model’s performance after densification closely matches that of the fully-trained model. The iterations after the densification stage essentially fine-tune the model and can be used to further “refine” it after pruning, similar to PUP 3D-GS [11] and LightGaussian [6]. In practice, our Hard Pruning method prunes the model by a constant ratio every $3000$ iterations starting at iteration $15000$. We Hard Prune $30\%$ of Gaussians in each interval, which, when paired with Soft Pruning, empirically reduces the total number of Gaussians across scenes by 10.6×10.6$\times$10.6 ×.
@@ 5.2 Implementation Details
Table 2: Average Gaussian count, FPS, and training time across all scenes in Section 5.1. Ratios for model size compression, rendering speed-up, and training speed-up are reported in (parentheses). The operation in each row is applied cumulatively to all of the following rows. The best and second best value for each metric are color coded.
@@
Table 3: Average reported metrics for each pruning method across all scenes in the Mip-NeRF 360 dataset. The Comp column reports model size compression in terms of Gaussian count, FPS reports rendering speed-up, and Train reports training time speed-up, all with respect to the baseline 3D-GS model. PSNR, SSIM, and LPIPS are also recorded. For a fair comparison, we report the published results of each method and use ‘-’ to denote missing metrics. The best and second best value for each metric are color coded; lossless methods are underlined. Results for the Tanks & Temples and Deep Blending datasets are reported in Appendix A.3.
@@ 5.3.2 Overall Performance
Table 3 compares our methods against other methods that reduce the number of Gaussians and increase inference speed using the mean of each metric across all scenes in the Mip-NeRF 360 dataset. The underlined methods are “lossless”, meaning that they avoid degrading visual fidelity at all. SnugBox and AccuTile, our lossless methods, improve rendering and training speed while leaving image quality metrics essentially unchanged or slightly better. Our full pipeline, labeled as “+Hard Pruning" boasts the highest compression ratios, rendering speeds, and training speed-ups across all datasets. Furthermore, its image quality metrics are competitive with the other methods.
@@ 5.4 Pruning Score Comparison
Table 4: Average metrics across all scenes in Section 5.1 when using the Speedy-Splat and PUP 3D-GS pruning scores to prune $88.44\%$ of Gaussians using the PUP 3D-GS pipeline. Two rounds of prune-refine are performed on each baseline 3D-GS model, pruning $66\%$ of Gaussians and then fine-tuning for $5,000$ iterations in each one. The best value for each metric is color coded.
@@ Appendix A Appendix
A.1 AccuTile Proof of Correctness Sketch
We outline the correctness of our AccuTile algorithm by examining the different cases that arise when identifying the minimum and maximum points of an ellipse within a given tile row. Due to the symmetry of Equation 14, exchanging the variables $x$ and $y$ along with the coefficients $a$ and $c$ yields equivalent statements for tile columns. Thus, we focus our discussion on tile rows.
Case 1: The ellipse does not intersect the tile row boundary. The entire ellipse, including the bounding box extrema $x_{\min}$ and $x_{\max}$ computed by SnugBox, lies within the row. AccuTile correctly selects these points as the furthest ellipse extents. Figure 4 illustrates an example of this.
Case 2: The ellipse intersects one of the tile row bounding lines but not the other.
This implies that either $y_{\min}$ or $y_{\max}$ lies within the row, but not both. There are several possible subcases:
  • Case 2.1: If both $x_{\min}$ and $x_{\max}$ are in the tile row, then they are correctly assigned as the furthest extent of the ellipse by AccuTile. Figure 5 illustrates an example of this case.
  • Case 2.2: If $x_{\min}$ and $x_{\max}$ are not in the row but $y_{\max}$ is, then the ellipse decreases monotonically from $y_{\max}$ to the row boundary on both sides of $y_{\max}$, making the row boundary intersections the furthest extent of the ellipse and are the points selected by AccuTile as the furthest row extent. This follows from the absence of the critical points $x_{\min}$ and $x_{\max}$. A symmetric argument applies when $y_{\min}$ is in the row instead. The top tile row of Figures 5 and 6 illustrate examples of this case.
  • Case 2.3: If $x_{\min}$ and $y_{\min}$ are in the row but $x_{\max}$ is not, then $x_{\min}$ is assigned as the minimum extent of the ellipse. The ellipse increases monotonically from $y_{\min}$ to the boundary to the right of $y_{\min}$ and from $x_{\min}$ to the boundary to the right of $x_{\min}$. Under a corrollary of the definition of an ellipse, the sides of the ellipse do not intersect. Thus, the ellipse point on the curve that extends to the right of $y_{\min}$ and intersects the tile row boundary must be the maximum ellipse extent, and AccuTile correctly selects it as such. A similar argument applies in the following cases: (1) $x_{\max}$ and $y_{\min}$ are in the tile row but $x_{\min}$ is not, (2) $x_{\min}$ and $y_{\max}$ are in the tile row but $x_{\max}$ is not, and (3) $x_{\max}$ and $y_{\max}$ are in the tile row but $x_{\min}$ is not. The bottom tile row of Figure 6 illustrates an example of this case.
Case 3: The ellipse intersects both the top and bottom row boundary.
If $x_{\min}$ or $x_{\max}$ is in the tile row, then AccuTile correctly assigns it as the minimum or maximum extent of the ellipse, respectively. The right side of the ellipse in the middle tile row in Figure 6 illustrates an example of this case. Otherwise, the ellipse monotonically increases from the bottom row boundary to the top row boundary, or vice-versa, due to the absence of critical points. Selecting the minimum or maximum boundary point, as done by AccuTile, yields the correct result. The left side of the ellipse in the middle tile row in Figure 6 illustrates an example of this case.
Figure 4: (Left) SnugBox and (right) AccuTile sketch of Case 1. As with Figure 2(c), our AccuTile algorithm iterates over the tile rows; the only points that are processed are $x_{min}$ and $x_{max}$.
Figure 5: (Left) SnugBox and (right) AccuTile sketch of Case 2.1. Our AccuTile algorithm iterates over the tile rows; the only points that are processed are $x_{min}$, $x_{max}$, A, and B.
Figure 6: (Left) SnugBox and (right) AccuTile sketch of Cases 2.2, 2.3, and 3. Our AccuTile algorithm iterates over the tile rows; the only points that are processed are $x_{min}$, $x_{max}$, A, B, C, and D. A detailed walkthrough of this example is presented in Section 4.1.2.
Figure 7: We sweep pruning percentages in $5\%$ increments for Hard Pruning ($0\%-40\%$) and Soft Pruning ($0\%$, $50\%-95\%$) on all scenes listed in Section 5.1. Experiments are performed 3×3$\times$3 × on each scene without our Gaussian localization methods; the reported metrics are averaged across all runs. ($0\%$$0\%$) is the baseline 3D-GS model, the first column ($0\%$, :) is Hard Pruning in isolation, and the first row (:, $0\%$) is Soft Pruning in isolation. The red dots at ($80\%$, $30\%$) denote the percentage settings used in our manuscript. We report the FPS increase and the Number of Gaussians and Train Time decrease factors to be consistent with the format in Table 3.
@@ A.1 AccuTile Proof of Correctness Sketch
A.2 Overall Pruning Percent Metrics
In Figure 7, we perform a parameter sweep over Hard Pruning percentages from $0\%-40\%$ at $5\%$ intervals and Soft Pruning percentages at $0\%$ and from $50-95\%$ at $5\%$ intervals. We conduct each experiment 3×3$\times$3 × on each scene listed in Section 5.1 to reduce variance, then average the metrics across all runs. All experiments are run without our Gaussian localization methods – SnugBox and AccuTile – to ablate the effect of each pruning method in isolation. Our ($80\%$, $30\%$) pruning percentages are empirically selected to produce a favorable balance between speed and quality.
@@ A.2 Overall Pruning Percent Metrics
A.3 Additional Datasets Evaluation
Table 7: Average execution time (milliseconds) of each function across all scenes. This experiment ablates the StopThePop [27] Tile-Based Culling method with warp-level load balancing against our AccuTile algorithm. For each method, execution times are averaged over three runs, with each run rendering the test set 20 times to reduce variance. The fastest times are highlighted. Our AccuTile algorithm outperforms the Tile-Based Culling method in overall runtime by a notable margin. For detailed analysis, see Section A.4.
Baseline 0.665 0.046 0.568 1.551 0.083 4.469 7.457
Tile-Based Culling [27] 0.811 (0.820x) 0.046 0.450 (1.263x) 0.609 (2.548x) 0.035 (2.341x) 2.027 (2.205x) 4.051 (1.841x)
AccuTile (Ours) 0.659 0.046 0.194 (2.931x) 0.610 (2.541x) 0.035 (2.338x) 2.042 (2.189x) 3.660 (2.038x)
Table 8: Average execution time (milliseconds) of each scene. This experiment ablates the StopThePop [27] Tile-Based Culling method with warp-level load balancing against our AccuTile algorithm by breaking out the per-scene execution times, which were averaged in the Overall column of Table 7. The fastest times are highlighted. Our AccuTile algorithm outperforms Tile-Based Culling on all scenes.
Mip-NeRF 360 Tanks & Temples Deep Blending
Baseline 14.034 4.977 7.025 7.914 6.103 11.025 8.562 5.776 7.088 6.993 4.872 7.253 5.322
Tile-Based Culling [27] 6.609 2.674 3.261 3.474 3.888 6.994 4.741 2.997 3.058 4.417 2.860 4.131 3.554
AccuTile (Ours) 5.880 2.401 2.989 2.933 3.478 6.433 4.490 2.578 2.726 3.965 2.729 3.671 3.300
A.4 StopThePop Tile-Based Culling Ablation
We ablate our AccuTile algorithm against the StopThePop [27] Tile-Based Culling method in Tables 7 and 8. Tile-Based Culling computes a precise Gaussian-to-Tile mapping in two steps: (1) Similar to our SnugBox method, a tight, opacity-aware bounding box is computed per Gaussian; however, due to the use of thresholds in their code, not all bounding boxes are tight. (2) Each tile touching the bounding box is iteratively examined to determine if it should be included in the final Gaussian-to-Tile mapping; warp-level load balancing is used to accelerate this process.
For this ablation, we update the 3D-GS rasterizer with the Tile-Based Culling code to isolate its runtime speed-up. All warp-level load balancing code is included to ensure that we compare against the most optimized version of the method. As noted in the StopThePop codebase, a padded alpha threshold is required to accurately compute bounding boxes, which, by extension, prevents undercounting Gaussian-to-Tile mappings. No padded alpha threshold is provided, so we perform this ablation without it. To ensure a fair comparison, we train three models for each scene and measure execution times with the baseline 3D-GS, Tile-Based Culling, and AccuTile renderers on each one.
Table 7 shows that our AccuTile method significantly outperforms Tile-Based Culling on Preprocess and Duplicate with Keys. Since Tile-Based Culling iterates over all candidate tiles while AccuTile does not, it requires more computation and induces a markedly higher runtime cost even with warp-level load-balancing. Surprisingly, Tile-Based Culling slightly outperforms AccuTile in the downstream functions Radix Sort, Identify Tile Ranges, and Render. However, we observe that this is caused by the aforementioned under-counting of Gaussian-to-Tile mappings; this marginal improvement disappears when padded alpha thresholds are introduced, further slowing down Preprocess and Duplicate with Keys. Additionally, as reported by Table 8, our AccuTile method consistently outperforms Tile-Based Culling across all scenes.
A.5 Per-Scene Metrics
PSNR, SSIM, LPIPS, FPS, and training times for each scene from the Mip-NeRF 360, Tanks&Temples, and Deep Blending datasets that was used in 3D-GS [14] are recorded in Tables 9, 10, 11,  12, and Table 13, respectively. The operation in each row is applied cumulatively to all of the following rows.